# Data-Driven Room Acoustic Modeling Via Differentiable Feedback Delay Networks With Learnable Delay Lines

### v2

@@ 6 Conclusions
We proposed a method for optimizing all the parameters of a time-invariant frequency-independent Feedback Delay Network (FDN) in order to match the reverberation of a given room through perceptually meaningful metrics. In the present work,
@@
In particular, we presented a new differentiable FDN that is characterized by learnable delay lines realized exploiting operations in the frequency domain. Thus, we jointly trained all FDN parameters via backpropagation taking into account a composite loss consisting of two terms: the normalized mean square error between target and predicted EDCs and the mean square error between target and predicted Soft EDPs. We tested the methodology on three RIRs of real rooms taken from a publicly available dataset, and we demonstrated that the Soft EDP term is essential for obtaining an IR that resembles a realistic RIR. Finally, we evaluated the methods against three baseline methods considering widespread metrics, including reverberation time, clarity, definition, and center time. Overall, the proposed approach was able to outperform the baselines across different metrics.
Future work includes the application of the proposed approach to frequency-dependent FDNs, which are able to account for a frequency-specific decay in time, or to multiple-input multiple-output (MIMO) delay networks.
@@
  • Authors’ contributions: A. I. Mezza conceptualized the study, designed and implemented the proposed method, run the experiments, and drafted the manuscript. R. Giampiccolo contributed to the design of the proposed method, run the experiments, and drafted part of the manuscript. E. De Sena contributed to the method and codebase, and drafted part of the manuscript. A. Bernardini revised the manuscript and supervised the work.

### v3

@@ 6 Conclusions
In this work, we proposed a method for optimizing every parameter of a time-invariant frequency-independent Feedback Delay Network (FDN) so as to match the reverberation of a given room through perceptually meaningful metrics. The main contributions are the following:
@@
In particular, we presented a new differentiable FDN that is characterized by learnable delay lines realized exploiting operations in the frequency domain. Thus, we jointly trained all FDN parameters via backpropagation taking into account a composite loss consisting of two terms: the normalized mean square error between target and predicted backward-integrated EDCs and the mean square error between target and predicted Soft EDPs. We evaluated the proposed method on three real-world RIRs taken from a publicly available dataset, and we demonstrated that the Soft EDP term is essential for obtaining an IR that resembles a realistic RIR. Finally, we tested our approach against three baseline methods considering widespread metrics, including reverberation time, clarity, definition, and center time. Overall, the proposed approach was able to outperform the baseline methods by a large margin across different metrics.
Future work includes the application of the proposed framework to frequency-dependent FDNs, which are able to account for a frequency-specific decay in time, or to multiple-input multiple-output (MIMO) delay networks.
@@
  • Authors’ contributions: A. I. Mezza conceptualized the study, designed and implemented the proposed method, run the experiments, and drafted the manuscript. R. Giampiccolo contributed to the design of the proposed method, run the experiments, and drafted the manuscript. E. De Sena contributed to the method and codebase and drafted part of the manuscript. A. Bernardini revised the manuscript and supervised the work.